### v1

# Bound States in Lee’s Complex Ghost Model

Abstract
Quantum field theories (QFTs) including fourth-derivative terms such as the Lee-Wick finite QED and quadratic gravity have a better ultra-violet behavior compared to standard theories with second-derivative ones, but the existence of ghost with negative norm endangers unitarity. Such a ghost in general acquires a pair of complex conjugate masses from radiative corrections whose features are concisely described by the so-called Lee model. Working with the canonical operator formalism of QFTs, we investigate the issue of bound states in the Lee model. We find that the bound states cannot be created from ghosts by contributions of a complex delta function, which is a complex generalization of the well-known Dirac delta function. Since the cause of unitarity violation in the Lee-Wick model is the existence of the complex delta function instead of the Dirac delta function, it is of interest to notice that the violation of the unitarity is also connected to the non-existence of bound states. Finally, the problem of amelioration of the unitarity in quadratic gravity is briefly discussed.
1 Introduction
It is well known that in higher derivative quantum field theories (QFTs), a ghost mode with negative metric (norm) appears in classical action either explicitly or effectively via radiative corrections. In particular, radiative corrections make the ghost mode with negative metric transform to a complex ghost. As for such a complex ghost, there is a surprising conjecture by Lee and Wick [1, 2] that the complex ghost cannot be created from scattering processes of physical particles with positive metric owing to energy conservation law of positive energy, and therefore the physical unitarity of the S-matrix is not violated.
However, this conjecture has been recently dismissed in Refs. [3, 4] where the complex delta function appearing at each vertex in Feyman diagrams plays an important role and leads to the violation of the physical unitarity of the S-matrix. In this respect, it is worthwhile to recall the difference between the physical unitarity and the total unitarity of the S-matrix: The physical unitarity of the S-matrix means that the probability interpretation of a theory is possible so that the presence of physical ghost particles carrying negative probability breaks this physical unitarity. On the other hand, the total unitarity of the S-matrix, i.e., $S^{\dagger}S=SS^{\dagger}=1$ is automatically obeyed as long as the Hamiltonian is Hermitian. The latter unitarity guarantees that the probability of a theory is conserved under time development regardless of whether there are normal particles with positive metric and ghost ones with negative metric.
More recently, based on path integral approach, using a rather simple scalar model which is very similar to the Lee’s complex ghost model, it is pointed out that a pair of complex ghosts forms a bound state with positive norm and thus leading to the confinement of ghost particles into a normal composite particle with positive norm [5, 6].
In this article, we wish to investigate a possibility that a pair of ghost particles is transformed to a normal bound state within the framework of the canonical operator formalism developed by Kubo and Kugo [3, 4]. Intuitively such the possibility would not be allowed by the conservation law of the total unitarity of the S-matrix by the following reasonings: Suppose that in the initial state there is a pair of ghost particles, each of which carries the value 1-1 of the ghost norm so totally we have 2-2 norm in the initial state. Then, whatever final states are produced from the initial state, the norm of the final states must sum up to the same value 2-2 as that of the initial state. Thus, this observation naturally leads to the conclusion that when only a single bound state is produced from a pair of ghost particles, it must possess the negative norm 2-2 and be a composite ghost. Of course, this observation does not exclude the existence of a bound state with positive norm if many of bound states and/or elementary fields are produced. The bottom line is that ghost particles and/or bound states with negative norm must be produced in the final states to match the value 2-2 after interactions.
The paper is organized as follows: In the next section, we introduce the Lee model and explain the canonical operator formalism of the model. The correlation function of a ghost composite operator is derived in Section 3 and the equation of pole positions is calculated in Section 4. Finally, in Section 5 we draw our conclusion.
2 Canonical operator formalism of Lee model
Let us start with a classical Lagrangian density of our model treated in this article:
${\cal{L}}={\cal{L}}^{(2)}+{\cal{L}}^{\rm{int}},$ (2.1)
where ${\cal{L}}^{(2)}$ and ${\cal{L}}^{\rm{int}}$ denote the quadratic kinetic term and the interaction one, respectively, whose concrete expressions read [^2] We choose the metric convention $\eta_{\mu\nu}=\rm{diag}(-1,+1,+1,+1)$.
${\cal{L}}^{(2)}$ $=$ $\frac{1}{2}\left[(\partial\varphi)^{2}+M^{2}\varphi^{2}+(\partial\varphi^{\dagger})^{2}+M^{*2}\varphi^{\dagger 2}\right],$
${\cal{L}}^{\rm{int}}$ $=$ $-\frac{f}{4}(\varphi^{\dagger}\varphi)^{2}.$ (2.2)
Here $\varphi$ is a complex scalar field and $\varphi^{\dagger}$ is its Hermitian conjugate, both of which denote complex ghost fields. $M^{2}$ denotes the complex squared mass of the scalar field and is assumed that both its real part ${\rm{Re}}\,M^{2}$ and imaginary one ${\rm{Im}}\,M^{2}$ are positive definite, i.e., ${\rm{Re}}\,M^{2}>0$ and ${\rm{Im}}\,M^{2}>0$. We will call the model defined by the above Lagrangian density the Lee model since although the original Lee model consists of a real scalar field of positive norm and that of negative norm together with their mixing term [7], we can diagonalize the Hamiltonian by introducing the complex ghost $\varphi$ and its Hermitian conjugate $\varphi^{\dagger}$ [8]. The imaginary part of $M^{2}$ may be an intrinstic one or may be due to radiative corrections as seen in a general Lagrangian density with fourth-derivative terms.
The free field equations for $\varphi$ and $\varphi^{\dagger}$ from ${\cal{L}}^{(2)}$ are
$(\Box-M^{2})\varphi=(\Box-M^{*2})\varphi^{\dagger}=0,$ (2.3)
so the complex scalar fields $\varphi$ and $\varphi^{\dagger}$ can be canonically quantized and are expanded as [8]
$\varphi(x)$ $=$ $\int\frac{d^{3}q}{\sqrt{(2\pi)^{3}2\omega_{q}}}\left[\alpha(\vec{q})e^{i\vec{q}\cdot\vec{x}-i\omega_{q}x^{0}}+\beta^{\dagger}(\vec{q})e^{-i\vec{q}\cdot\vec{x}+i\omega_{q}x^{0}}\right],$
$\varphi^{\dagger}(x)$ $=$ $\int\frac{d^{3}q}{\sqrt{(2\pi)^{3}2\omega_{q}^{*}}}\left[\alpha^{\dagger}(\vec{q})e^{-i\vec{q}\cdot\vec{x}+i\omega_{q}^{*}x^{0}}+\beta(\vec{q})e^{i\vec{q}\cdot\vec{x}-i\omega_{q}^{*}x^{0}}\right],$ (2.4)
where
$q^{\mu}\equiv(\omega_{q},\vec{q}),\qquad\omega_{q}\equiv\sqrt{\vec{q}\,^{2}+M^{2}},\qquad\omega_{q}^{*}\equiv\sqrt{\vec{q}\,^{2}+M^{*2}}.$ (2.5)
As a feature of the system with the complex mass, creation operators $\alpha^{\dagger}(\vec{q}),\beta^{\dagger}(\vec{q})$ and annihilation ones $\alpha(\vec{q}),\beta(\vec{q})$ obey the off-diagonal commutation relations [8]:
$[\alpha(\vec{p}),\beta^{\dagger}(\vec{q})]=[\beta(\vec{p}),\alpha^{\dagger}(\vec{q})]=-\delta^{3}(p-q),$
$[\alpha(\vec{p}),\alpha^{\dagger}(\vec{q})]=[\beta(\vec{p}),\beta^{\dagger}(\vec{q})]=0.$ (2.6)
Next, we turn our attention to the propagators of $\varphi$ and $\varphi^{\dagger}$. By means of Eqs. (2.4), (2.6) and the definition of the T-product, it is easy to obtain the propagator of $\varphi$ as follows [3, 4]:
$D_{\varphi}(x-y)\equiv\langle 0|T\varphi(x)\varphi(y)|0\rangle$
$\equiv\langle 0|[\theta(x^{0}-y^{0})\varphi(x)\varphi(y)+\theta(y^{0}-x^{0})\varphi(y)\varphi(x)]|0\rangle$
$=\int\frac{d^{3}qd^{3}q^{\prime}}{(2\pi)^{3}\sqrt{2\omega_{q}\cdot 2\omega_{q^{\prime}}}}\Big[\theta(x^{0}-y^{0})e^{i(\vec{q}\cdot\vec{x}-\omega_{q}x^{0}-\vec{q}\,^{\prime}\cdot\vec{y}+\omega_{q^{\prime}}y^{0})}\langle 0|\alpha(\vec{q})\beta^{\dagger}(\vec{q}\,^{\prime})|0\rangle$
$+\theta(y^{0}-x^{0})e^{i(\vec{q}\,^{\prime}\cdot\vec{y}-\omega_{q^{\prime}}y^{0}-\vec{q}\cdot\vec{x}+\omega_{q}x^{0})}\langle 0|\alpha(\vec{q}\,^{\prime})\beta^{\dagger}(\vec{q})|0\rangle\Big]$
$=-\int\frac{d^{3}q}{(2\pi)^{3}2\omega_{q}}\Big[\theta(x^{0}-y^{0})e^{i\vec{q}\cdot(\vec{x}-\vec{y})-i\omega_{q}(x^{0}-y^{0})}+\theta(y^{0}-x^{0})e^{-i\vec{q}\cdot(\vec{x}-\vec{y})+i\omega_{q}(x^{0}-y^{0})}\Big]$
$=-\int\frac{d^{3}q}{(2\pi)^{3}}e^{i\vec{q}\cdot(\vec{x}-\vec{y})}\int_{C}\frac{dq^{0}}{2\pi i}\frac{e^{-iq^{0}(x^{0}-y^{0})}}{q^{2}+M^{2}}$
$\equiv\int_{C}\frac{d^{4}q}{i(2\pi)^{4}}e^{iq\cdot(x-y)}D_{\varphi}(q),$ (2.7)
where we have defined
$D_{\varphi}(q)\equiv-\frac{1}{q^{2}+M^{2}}.$ (2.8)
Here the contour $C$ is the well-known Lee-Wick complex contour [7] which starts at $q^{0}=$-$\infty$, passes below the left pole at $q^{0}=-\omega_{q}$ and above the right pole at $q^{0}=\omega_{q}$, and then goes to the infinity $q^{0}=+\infty$ in the complex $q^{0}$-plane, as shown in Fig. 1. Note that this contour is automatically derived in the canonical formalism under consideration without recourse to the definition of the S-matrix [8].
$\mathrm{Re}\,q^{0}$$\mathrm{Im}\,q^{0}$$\omega_{\mathbf{q}}$$\omega_{\mathbf{q}}$0CCFig.1. The Lee-Wick contour
In a perfectly similar manner, the propagator of $\varphi^{\dagger}$ is obtained to be
$D_{\varphi^{\dagger}}(x-y)\equiv\langle 0|T\varphi^{\dagger}(x)\varphi^{\dagger}(y)|0\rangle$
$=-\int\frac{d^{3}q}{(2\pi)^{3}2\omega_{q}^{*}}\Big[\theta(x^{0}-y^{0})e^{i\vec{q}\cdot(\vec{x}-\vec{y})-i\omega_{q}^{*}(x^{0}-y^{0})}+\theta(y^{0}-x^{0})e^{-i\vec{q}\cdot(\vec{x}-\vec{y})+i\omega_{q}^{*}(x^{0}-y^{0})}\Big]$
$=-\int\frac{d^{3}q}{(2\pi)^{3}}e^{i\vec{q}\cdot(\vec{x}-\vec{y})}\int_{R}\frac{dq^{0}}{2\pi i}\frac{e^{-iq^{0}(x^{0}-y^{0})}}{q^{2}+M^{*2}}$
$\equiv\int_{R}\frac{d^{4}q}{i(2\pi)^{4}}e^{iq\cdot(x-y)}D_{\varphi^{\dagger}}(q),$ (2.9)
where we have also defined
$D_{\varphi^{\dagger}}(q)\equiv-\frac{1}{q^{2}+M^{*2}}.$ (2.10)
Note that in this case the contour must be chosen along the real axis $R$ because of $M^{*2}={\rm{Re}}\,M^{2}-i{\rm{Im}}\,M^{2}$ with ${\rm{Im}}\,M^{2}>0$.
3 Correlation function of composite operator
In this section, we will derive the correlation function of a composite operator of ghost fields on the basis of the canonical operator formalism of the Lee model as explained in the previous section. This problem has been recently discussed in Refs. [5, 6] from path integral approach. We will obtain the different expression of the correlation function from that of Refs. [5, 6]. It will turn out that the difference stems from a careful treatment of the Lee-Wick complex contour and the complex delta function.
We would like to consider a bound state problem constructed out of the following composite ghost operator:
${\cal{O}}_{|\varphi|^{2}}(x)\equiv\varphi^{\dagger}(x)\varphi(x).$ (3.1)
This is a neutral operator and is one of the most plausible operators if a bound state is made from ghost fields. Since the other ghost composite operators are dealed with in the same way, we will focus our attention on whether bound states are made from the operator ${\cal{O}}_{|\varphi|^{2}}(x)$ in Eq. (3.1) or not within the framework of the canonical operator formalism.
According to the Gell-Mann-Low formula [9], in the canonical operator formalism the correlation function of the ghost composite operator ${\cal{O}}_{|\varphi|^{2}}(x)$ is defined as [^3] See Refs. [10, 11, 12] for composite field operators.
$\langle{\cal{O}}_{|\varphi|^{2}}(x){\cal{O}}_{|\varphi|^{2}}(y)\rangle\equiv\langle{\bf{0}}|T[{\cal{O}}_{|\varphi|^{2}}(x){\cal{O}}_{|\varphi|^{2}}(y)]|{\bf{0}}\rangle$ (3.2)
$\equiv$ $\sum_{n=0}^{\infty}\frac{i^{n}}{n!}\lim_{\varepsilon\rightarrow+0}\int d^{4}x_{1}\cdots d^{4}x_{n}\,{\rm{exp}}(-\varepsilon\sum_{i=1}^{n}|x_{i}^{0}|)\frac{1}{\langle 0|S|0\rangle}$
$\times$ $\langle 0|T[{\cal{L}}^{\rm{int}}(x_{1})\cdots{\cal{L}}^{\rm{int}}(x_{n})\varphi^{\dagger}(x)\varphi(x)\varphi^{\dagger}(y)\varphi(y)]|0\rangle,$
[…]
$\langle{\cal{O}}_{|\varphi|^{2}}(x){\cal{O}}_{|\varphi|^{2}}(y)\rangle$ (3.3)
$=$ $D_{\varphi^{\dagger}}(x-y)D_{\varphi}(x-y)-if\int d^{4}z\,D_{\varphi^{\dagger}}(x-z)D_{\varphi^{\dagger}}(z-y)D_{\varphi}(x-z)D_{\varphi}(z-y)$
- $f^{2}\int d^{4}zd^{4}w\,D_{\varphi^{\dagger}}(x-z)D_{\varphi^{\dagger}}(z-w)D_{\varphi^{\dagger}}(w-y)D_{\varphi}(x-z)D_{\varphi}(z-w)D_{\varphi}(w-y)$
$+$ ${\cal{O}}(f^{3}).$
From now on, let us explicitly calculate each term on the right hand side (RHS) of this equation. As for the first term, $D_{\varphi^{\dagger}}(x-y)D_{\varphi}(x-y)$, paying an attention to the difference of the contours $C$ and $R$, the calculation proceeds as follows:
$D_{\varphi^{\dagger}}(x-y)D_{\varphi}(x-y)$ (3.4)
$=$ $\int_{R}\frac{d^{4}q}{i(2\pi)^{4}}e^{iq\cdot(x-y)}D_{\varphi^{\dagger}}(q)\int_{C}\frac{d^{4}q^{\prime}}{i(2\pi)^{4}}e^{iq^{\prime}\cdot(x-y)}D_{\varphi}(q^{\prime})$
$=$ $\int_{C}\frac{d^{4}k}{i(2\pi)^{4}}e^{ik\cdot(x-y)}\int_{R}\frac{d^{4}q}{i(2\pi)^{4}}D_{\varphi^{\dagger}}(q)D_{\varphi}(k-q)$
$\equiv$ $\int_{C}\frac{d^{4}k}{i(2\pi)^{4}}e^{ik\cdot(x-y)}G(k),$
where $G(k)$ is defined by
$G(k)\equiv\int_{R}\frac{d^{4}q}{i(2\pi)^{4}}D_{\varphi^{\dagger}}(q)D_{\varphi}(k-q).$ (3.5)
Note that at the second equality of Eq. (3.4) we have made the change of variable from $q^{\prime}$ to $k$, which is defined as $k=q+q^{\prime}$, but $q^{0}$ and $q^{\prime\,0}$ run along the contours $R$ and $C$, respectively, so the variable $q^{\prime}$, not $q$, must be expressed in terms of $k$.
Next, let us move on to the second term on the RHS of Eq. (3.3). Up to an overall factor if-if, this term can be evaluated to be
$\int d^{4}z\,D_{\varphi^{\dagger}}(x-z)D_{\varphi^{\dagger}}(z-y)D_{\varphi}(x-z)D_{\varphi}(z-y)$ (3.6)
$=$ $\int d^{4}z\,D_{\varphi^{\dagger}}(x-z)D_{\varphi}(x-z)D_{\varphi^{\dagger}}(z-y)D_{\varphi}(z-y)$
$=$ $\int d^{4}z\,\int_{C}\frac{d^{4}k}{i(2\pi)^{4}}\int_{C}\frac{d^{4}k^{\prime}}{i(2\pi)^{4}}e^{i(kx-k^{\prime}y)}e^{-i(k-k^{\prime})z}G(k)G(k^{\prime}).$
[…]
$\frac{1}{(2\pi)^{4}}\int d^{4}z\,e^{-i(k-k^{\prime})z}=\delta_{c}(k^{0}-k^{\prime\,0})\delta^{3}(k-k^{\prime}),$ (3.7)
where the complex delta function is defined as
$\frac{1}{2\pi}\int dz^{0}\,e^{-i(k^{0}-k^{\prime\,0})z^{0}}=\delta_{c}(k^{0}-k^{\prime\,0}),$ (3.8)
for general complex numbers $k^{0}$ and $k^{\prime\,0}$. Thus we arrive at
$\int d^{4}z\,D_{\varphi^{\dagger}}(x-z)D_{\varphi^{\dagger}}(z-y)D_{\varphi}(x-z)D_{\varphi}(z-y)$ (3.9)
$=$ $\int_{C}\frac{d^{4}kdk^{\prime\,0}}{i^{2}(2\pi)^{4}}e^{-i(k^{0}x^{0}-k^{\prime\,0}y^{0})+i\vec{k}\cdot(\vec{x}-\vec{y})}\delta_{c}(k^{0}-k^{\prime\,0})G(k^{0},\vec{k})G(k^{\prime\,0},\vec{k}).$
In a perfectly similar manner, one can calculate the third term on the RHS of Eq. (3.3). Putting all terms together, we have an expression for the correlation function:
$\langle{\cal{O}}_{|\varphi|^{2}}(x){\cal{O}}_{|\varphi|^{2}}(y)\rangle\equiv\int_{C}\frac{d^{4}k}{i(2\pi)^{4}}\,e^{ik\cdot(x-y)}C(k),$ (3.10)
where $C(k)$ is defined by
$C(k)$ $\equiv$ $G(k)\Big[1-f\int_{C}dk^{\prime\,0}\,\delta_{c}(k^{0}-k^{\prime\,0})G(k^{\prime\,0},\vec{k})$ (3.11)
$+$ $f^{2}\int_{C}dk^{\prime\,0}dk^{\prime\prime\,0}\,\delta_{c}(k^{0}-k^{\prime\,0})\delta_{c}(k^{0}-k^{\prime\prime\,0})G(k^{\prime\,0},\vec{k})G(k^{\prime\prime\,0},\vec{k})$
$+$ ${\cal{O}}(f^{3})\Big],$
where we have used the following identity:
[…]
which can be easily proved.
Moreover, we can rewrite $C(k)$ into a more concise form:
$C(p)$ $=$ $\frac{G(p)}{1+f\int_{C}dp^{\prime\,0}\,\delta_{c}(p^{0}-p^{\prime\,0})G(p^{\prime\,0},\vec{p})}$
$=$ $\frac{G(p)}{1+f\int_{C}d^{4}k\int_{R}\frac{d^{4}q}{i(2\pi)^{4}}\,D_{\varphi}(k)D_{\varphi^{\dagger}}(q)\delta_{c}(k^{0}+q^{0}-p^{0})\delta^{3}(k+q-p)},$
where we have used Eq. (3.5) in the last line. This equation implies that if there were a pole at some on-shell value in Eq. (LABEL:C(p)), $p^{2}=-m^{2}$, we would have a bound state with the mass $m$. Thus, by solving this equation in the next section, we will investigate the problem of whether there is such a bound state or not.
4 Bound states
As explained in the previous section, when the pole equation
$1+f\int_{C}d^{4}k\int_{R}\frac{d^{4}q}{i(2\pi)^{4}}\,D_{\varphi}(k)D_{\varphi^{\dagger}}(q)\delta_{c}(k^{0}+q^{0}-p^{0})\delta^{3}(k+q-p)=0,$ (4.1)
has a solution at the on-shell value, $p^{2}=-m^{2}$, one can judge that there is a bound state with the mass $m$. Furthermore, when the bound state has a positive residue at the pole, it is a physical bound state whereas when the bound state has a negative residue, it is a unphysical ghost bound state. Anyway, we have to verify that there is a solution to the pole equation (4.1).
For this aim, let us evaluate
$J\equiv\int_{C}d^{4}k\int_{R}d^{4}q\,D_{\varphi}(k)D_{\varphi^{\dagger}}(q)\delta_{c}(k^{0}+q^{0}-p^{0})\delta^{3}(k+q-p).$ (4.2)
Then, the pole equation (4.1) is simply written as
$1+f\frac{1}{i(2\pi)^{4}}J=0.$ (4.3)
After integrating over $\vec{q}$, $J$ is cast to the form:
$J=\int d^{3}k\int_{C}dk^{0}\int_{R}dq^{0}\,\frac{1}{-(k^{0})^{2}+\omega_{k}^{2}}\frac{1}{-(q^{0})^{2}+(\omega_{p-k}^{*})^{2}}\delta_{c}(k^{0}+q^{0}-p^{0}),$ (4.4)
where
$\omega_{k}\equiv\sqrt{\vec{k}^{2}+M^{2}},\qquad\omega_{p-k}^{*}\equiv\sqrt{(\vec{p}-\vec{k})^{2}+M^{*2}}.$ (4.5)
In evaluating this integral, the point is that $k^{0}$ is in general complex and takes values along the greatly deformed Lee-Wick contour $C$ in the $k^{0}$-plane, as seen in Fig. 1. This $k^{0}$-integration along the contour $C$ is then reduced to the sum of the integration over the real axis $R$ and residues at two poles, $k^{0}=\pm\omega_{k}$:
$\int_{C}dk^{0}\rightarrow\left(\int_{R}+\int_{C(-\omega_{k})}-\int_{C(\omega_{k})}\right)dk^{0},$ (4.6)
where $C(\omega)$ denotes an infinitesimal circle rotating anti-clockwise around the pole at $k^{0}=\omega$, and the minus sign in front of the last term on the RHS comes from the clockwise rotation. Consequently, we can obtain
$J=\int d^{3}k\int_{R}dq^{0}\Bigg[\int_{R}dk^{0}\frac{1}{-(k^{0})^{2}+\omega_{k}^{2}}\delta(k^{0}+q^{0}-p^{0})+\frac{\pi i}{\omega_{k}}\sum_{\pm}\delta_{c}(q^{0}\pm\omega_{k}-p^{0})\Bigg]$ (4.7)
$\times$ $\frac{1}{-(q^{0})^{2}+(\omega_{p-k}^{*})^{2}}$
$=$ $\int d^{3}k\Bigg[\int_{R}dk^{0}\frac{1}{-(k^{0})^{2}+\omega_{k}^{2}}\frac{1}{-(k^{0}-p^{0})^{2}+(\omega_{p-k}^{*})^{2}}+\frac{\pi i}{\omega_{k}}\sum_{\pm}\int_{R}dq^{0}\delta_{c}(q^{0}\pm\omega_{k}-p^{0})$
$\times$ $\frac{1}{-(q^{0})^{2}+(\omega_{p-k}^{*})^{2}}\Bigg],$
where we have defined
$\sum_{\pm}\delta_{c}(q^{0}\pm\omega_{k}-p^{0})=\delta_{c}(q^{0}+\omega_{k}-p^{0})+\delta_{c}(q^{0}-\omega_{k}-p^{0}),$ (4.8)
and used the fact that $\delta_{c}(k^{0}+q^{0}-p^{0})\rightarrow\delta(k^{0}+q^{0}-p^{0})$ in the first line of Eq. (4.7) since $k^{0},q^{0}$ and $p^{0}$ are all real numbers in this case.
The last term in the second line of Eq. (4.7) can be simplified by taking account of a suitable rectangular region whose form depends on $\delta_{c}(q^{0}-\omega_{k}-p^{0})$ or $\delta_{c}(q^{0}+\omega_{k}-p^{0})$. In case of $\delta_{c}(q^{0}-\omega_{k}-p^{0})$, the rectangular region is surrounded by the contour which is consisted of the real axis $R$, the finite vertical segment $C_{1}$ located at $q^{0}=$-$\infty$, the horizontal line $R(\omega_{k})$ parallel to the real axis and running from $\infty+i{\rm{Im}}\,\omega_{k}$ to $\infty+i{\rm{Im}}\,\omega_{k}$ and the finite vertical segment $C_{2}$ located at $q^{0}=+\infty$. Furthermore, we take this contour in order not to include the infinitesimal circle rotating anti-clockwise around the pole at $q^{0}=-\omega_{p-k}^{*}$.
[…]
Since there is no pole inside such a rectangular region, we have
$\int_{R}dq^{0}\delta_{c}(q^{0}-\omega_{k}-p^{0})\frac{1}{-(q^{0})^{2}+(\omega_{p-k}^{*})^{2}}$ (4.9)
$=$ $\left(\int_{R(\omega_{k})}+\int_{C(-\omega_{p-k}^{*})}\right)dq^{0}\delta_{c}(q^{0}-\omega_{k}-p^{0})\frac{1}{-(q^{0})^{2}+(\omega_{p-k}^{*})^{2}},$
where we have used that the integrals along the finite vertical segments $C_{1}$ and $C_{2}$ vanish. The first integral along $R(\omega_{k})$ is easily evaluated by making the change of variable from $q^{0}$ to $q^{\prime\,0}\equiv q^{0}-i\,{\rm{Im}}\,\omega_{k}$, by which the complex delta function is reduced to the Dirac delta function, and then the second integral is done by calculating the residue at the pole $q^{0}=-\omega_{p-k}^{*}$.
$\int_{R}dq^{0}\delta_{c}(q^{0}-\omega_{k}-p^{0})\frac{1}{-(q^{0})^{2}+(\omega_{p-k}^{*})^{2}}$ (4.10)
$=$ $\frac{1}{-(p^{0}+\omega_{k})^{2}+(\omega_{p-k}^{*})^{2}}+\frac{\pi i}{\omega_{p-k}^{*}}\delta_{c}[-(\omega_{p-k}^{*}+\omega_{k})-p^{0}].$
In a perfectly similar way, one can also calculate the case where $\delta_{c}(q^{0}-\omega_{k}-p^{0})$ is replaced with $\delta_{c}(q^{0}+\omega_{k}-p^{0})$. The two results are summarized in
$\int_{R}dq^{0}\delta_{c}(q^{0}\pm\omega_{k}-p^{0})\frac{1}{-(q^{0})^{2}+(\omega_{p-k}^{*})^{2}}$ (4.11)
$=$ $\frac{1}{-(p^{0}\mp\omega_{k})^{2}+(\omega_{p-k}^{*})^{2}}+\frac{\pi i}{\omega_{p-k}^{*}}\delta_{c}[\pm(\omega_{p-k}^{*}+\omega_{k})-p^{0}].$
Accordingly, $J$ reads
$J=\int d^{3}k\Bigg\{\int_{R}dk^{0}\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}+\frac{\pi i}{\omega_{k}}\sum_{\pm}\frac{1}{-(p^{0}\mp\omega_{k})^{2}+(\omega_{p-k}^{*})^{2}}$ (4.12)
- $\frac{\pi^{2}}{\omega_{k}\omega_{p-k}^{*}}\sum_{\pm}\delta_{c}[\pm(\omega_{p-k}^{*}+\omega_{k})-p^{0}]\Bigg\},$
Incidentally, it turns out that the first and second terms can be cast to the integral along the Lee-Wick contour $C$:
$J=\int_{C}d^{4}k\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}-\int d^{3}k\frac{\pi^{2}}{\omega_{k}\omega_{p-k}^{*}}\sum_{\pm}\delta_{c}[\pm(\omega_{p-k}^{*}+\omega_{k})-p^{0}].$
Now we are ready to calculate the first term on the RHS of Eq. (4.12). This integral is a usual one-loop integral except for complex masses. First, we would like to make the Wick rotation, by which the $k^{0}$ integration contour along the real axis is rotated to the imaginary axis. In contrast to the conventional situation, the contour now passes over two poles at $k^{0}=\pm\omega_{k}$. [^6] In addition to the two poles, the contour might pass over a pole at $k^{0}=\sqrt{\vec{p}\,^{2}+m^{2}}-\omega_{p-k}^{*}$. Here we assume ${\rm{Re}}(\sqrt{\vec{p}\,^{2}+m^{2}}-\omega_{p-k}^{*})<0$, which depends on the magnitudes of $m,{\rm{Re}}\,M^{2}$ and $|\vec{k}|$. If not, we would get an additional residue at this pole, but our final conclusion remains unchanged. The evaluation of contributions from the poles leads to
$\int_{R}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}$
$=$ $\int_{I}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}-\int d^{3}k\,\frac{\pi i}{\omega_{k}}\sum_{\pm}\frac{1}{-(p^{0}\mp\omega_{k})^{2}+(\omega_{p-k}^{*})^{2}},$
where $I$ denotes the imaginary axis running from $(0,-i\infty)$ to $(0,i\infty)$ in the complex $k^{0}$-plane. Note that the second term just cancels the second term of Eq. (4.12). To put it differently, we can rewrite the first and second terms of Eq. (4.12) into the integral over the imaginary axis [^7] In Ref. [5], the authors start with the Euclidean Lagrangian density and calculate only the same quantity as $K(p)$ on the basis of path integral approach.
$K(p)$ $\equiv$ $\int_{R}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}+\int d^{3}k\,\frac{\pi i}{\omega_{k}}\sum_{\pm}\frac{1}{-(p^{0}\mp\omega_{k})^{2}+(\omega_{p-k}^{*})^{2}}$ (4.15)
$=$ $\int_{I}d^{4}k\,\frac{1}{k^{2}+M^{2}}\frac{1}{(p-k)^{2}+M^{*2}}.$
Next it is straightforward to calculate $K$. We introduce $k^{4}$ defined by $k^{0}=ik^{4}$. Then we have $k^{2}=-(k^{0})^{2}+\vec{k}^{2}=(k^{4})^{2}+\vec{k}^{2}\equiv k_{E}^{2}$. Using these Euclidean momenta, $K(p)$ can be recast to the form
$K(p)=i\int d^{4}k_{E}\,\frac{1}{k_{E}^{2}+M^{2}}\frac{1}{(p_{E}-k_{E})^{2}+M^{*2}},$ (4.16)
where we have introduced the Euclidean momenta $p_{E}$ as well. Below we will write $k,p$ instead of $k_{E},p_{E}$ for convenience. By means of the Feyman’s parameter formula, $K(p)$ can be rewritten as follows:
$K(p)$ $=$ $i\int d^{4}k\int_{0}^{1}dx\,\frac{1}{\{[k-p(1-x)]^{2}+p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x\}^{2}}$ (4.17)
$=$ $i\int d^{4}k\int_{0}^{1}dx\,\frac{1}{[k^{2}+p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x]^{2}}$
$=$ $i\int_{0}^{1}dx\int d^{4}k\,\frac{1}{(k^{2}+\Delta)^{2}},$
where in the second line we have shifted the integration variable [^8] Precisely speaking, this procedure is legitimate only after regularizing the integral in a proper manner. as $k^{\mu}\rightarrow k^{\mu}+p^{\mu}(1-x)$ and in the last line we have defined
$\Delta\equiv p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x.$ (4.18)
It is obvious that $K(p)$ is logarithmically divergent so we adopt the Pauli-Villars regularization. To do that, let us consider the following quantity:
$I(p)=\int d^{4}k\,\left[\frac{1}{(k^{2}+\Delta)^{2}}-\frac{1}{(k^{2}+\Lambda^{2})^{2}}\right].$ (4.19)
Then, $I(p)$ can be calculated to be
$I(p)$ $=$ $\int d^{4}k\int_{\Delta}^{\Lambda^{2}}dy\,\frac{2}{(k^{2}+y)^{3}}$ (4.20)
$=$ $4\pi^{2}\int_{\Delta}^{\Lambda^{2}}dy\int_{0}^{\infty}dk\,\frac{k^{3}}{(k^{2}+y)^{3}}$
$=$ $-\pi^{2}\log\frac{\Delta}{\Lambda^{2}},$
where we have used the spherical coordinates in the second line. Thus, from Eqs. (4.17)-(4.20), $K(p)$ takes the form:
$K(p)$ $=$ $-i\pi^{2}\int_{0}^{1}dx\,\log\frac{p^{2}x(1-x)+M^{*2}+(M^{2}-M^{*2})x}{\Lambda^{2}}$ (4.21)
$=$ $-i\pi^{2}\Bigg\{\log\frac{-MM^{*}}{\Lambda^{2}}+\frac{M^{2}-M^{*2}}{p^{2}}\log\frac{M^{*}}{M}+\frac{a(p)}{p^{2}}$
$\times$ $\log\frac{[p^{2}-M^{2}+M^{*2}+a(p)][p^{2}+M^{2}-M^{*2}+a(p)]}{4MM^{*}p^{2}}\Bigg\},$
where
$a(p)\equiv\sqrt{(p^{2}+M^{2}-M^{*2})^{2}+4M^{*2}p^{2}}.$ (4.22)
Consequently, the pole equation (4.3), or equivalently (4.1), reads
$1+f\frac{1}{i(2\pi)^{4}}\Bigg(-i\pi^{2}\Bigg\{\log\frac{-MM^{*}}{\Lambda^{2}}+\frac{M^{2}-M^{*2}}{p^{2}}\log\frac{M^{*}}{M}+\frac{a(p)}{p^{2}}$ (4.23)
$\times$ $\log\frac{[p^{2}-M^{2}+M^{*2}+a(p)][p^{2}+M^{2}-M^{*2}+a(p)]}{4MM^{*}p^{2}}\Bigg\}$
- $\int d^{3}k\sum_{\pm}\frac{\pi^{2}}{\omega_{k}\omega_{p-k}^{*}}\delta_{c}[\pm(\omega_{p-k}^{*}+\omega_{k})-p^{0}]\Bigg)=0.$
Here let us suppose that there is no the last term on the LHS, which involves the complex delta function. Then, since the second term on the LHS including the logarithmic function, i.e., the function $K(p)$, has a rather smooth expression with respect to $p^{2}(=-m^{2})$, there could be a possibility such that we have a nontrivial solution to Eq. (4.23) as suggested in Ref. [5, 6]. However, the presence of the last term forbits such a solution. In other words, there is no non-trivial solution to Eq. (4.23) owing to the existence of the last term involving the complex delta function. It is worthwhile to stress that it is the term that breaks the unitarity of the Lee model [3, 4]. To put it differently, the term with the complex delta function not only breaks the unitarity but also prohibits the existence of the bound state made out of complex ghost fields.
Finally, let us comment on the other choices of composite operators constructed from ghost fields such as ${\cal{O}}(x)=\varphi^{\dagger 2}(x)+\varphi^{2}(x)$ etc. We have investigated several candidates of composite operators and found that the complex delta functions always appear and prohibit the existence of nontrivial solutions to the pole equation. It is therefore concluded that there is no bound state constructed from ghost fields in the Lee model.
5 Conclusions
In this paper, based on the canonical operator formalism we have demonstrated that there is no bound state in the Lee model which nicely describes some features of QFTs with fourth-derivative terms. We should stress that we have only made use of the operator formalism, which is the basic tool of QFTs.
Here it is interesting to imagine whether if there were a bound state with positive norm, it could give us a resolution to the problem of massive ghost, which violates the unitarity, in the higher derivative theories [14, 15, 16, 17, 18, 19, 20]. We can show that that is not the case from two reasons. One of the two reasons is that as mentioned in the introduction, even if there would be a bound state made from ghost fields, the bound state must be not a normal bound state with positive norm but a ghost bound state with negative norm by the conservation law of the total unitarity of the S-matrix. The other reason is that the construction of the bound state does not mean the permanent confinement of the massive ghost and the bound state is dissolved into elementary ghost fields in the weak coupling regime. To solve the problem of the massive ghost in the higher derivative theories such as quadratic gravity, we need the permanent confinement such as the confinement of quarks and gluons in QCD.
Under such a situation, it is valuable to pursue a new solution to the massive ghost problem from the viewpoint of bound states within the framework of the canonical formalism. For instance, in the analysis of physical modes of quadratic gravity [21] [^9] The BRST formalism of various gravitational theories has been already constructed in the de Donder (harmonic) gauge [22, 23, 24, 25, 26, 27, 28]., we usually assume that there is no bound state. If this assumption is not imposed, we are free to take account of a bound state which might be constructed from Faddeev-Popov (FP) ghosts and the other elementary fields. Of course, as mentioned above, such a bound state carries a negative norm so it must be a ghost bound state. However, if a certain bound state is constructed from the BRST ghost and its BRST-conjugate is also formed, they together constitute a BRST doublet, thereby belonging to unphysical state and nullifying the massive ghost problem [29] since this mechanism realizes the scenario of the permanent confinement of the massive ghost. The present formalism would be useful in clarifying the (non-)existence of such the bound state. We wish to return to this problem in the near future.